Henry Company's Accounts Receivable increased by $800 during the year. If they also reported Net Credit Sales of

$29,800, how much cash was received from their credit customers?

A)

$26,400.

B)

$29,000.

C)

$29,800.

D)

$30,600.

E)

$32,400.
Answer #1

Let beginning AR=$1000

Ending AR=($1000+800)=$1800

Ending AR=Beginning AR+Net Credit Sales-Cash collections

1800=1000+29800-Cash collections

Cash collections=1000+29800-1800

=$29000.